# Scalable Approximate Biclique Counting over Large Bipartite Graphs

### v1

@@ Abstract.
Counting $(p,q)$-bicliques in bipartite graphs is crucial for a variety of applications, from recommendation systems to cohesive subgraph analysis. Yet, it remains computationally challenging due to the combinatorial explosion to exactly count the $(p,q)$-bicliques. In many scenarios, e.g., graph kernel methods, however, exact counts are not strictly required. To design a scalable and high-quality approximate solution, we novelly resort to $(p,q)$-broom, a special spanning tree of the $(p,q)$-biclique, which can be counted via graph coloring and efficient dynamic programming. Based on the intermediate results of the dynamic programming, we propose an efficient sampling algorithm to derive the approximate $(p,q)$-biclique count from the $(p,q)$-broom counts. Theoretically, our method offers unbiased estimates with provable error guarantees. Empirically, our solution outperforms existing approximation techniques in both accuracy (up to 8$\times$ error reduction) and runtime (up to 50$\times$ speedup) on nine real-world bipartite networks, providing a scalable solution for large-scale $(p,q)$-biclique counting.
copyright: noneconference: ; ;
@@ 1. Introduction
The bipartite graph stands as a cornerstone in graph mining, comprising two distinct sets of vertices where edges exclusively link vertices from different sets. This concept serves as a powerful tool for modeling relationships across various real-world domains, including recommendation networks (Chen et al., 2020), collaboration networks (Beutel et al., 2013), and gene coexpression networks (Kaytoue et al., 2011). For example, in recommendation networks, users and items typically constitute two distinct vertex types. The interactions between these vertices form a bipartite graph, with edges representing users’ historical purchase behaviors.
@@
Despite its importance, counting $(p,q)$-bicliques is very challenging due to its exponential increase with respect to $p$ and $q$ (Yang et al., 2021). For example, in one of the graph datasets Twitter, with fewer than $2\times 10^{6}$ edges, there are more than $10^{13}$ $(5,4)$-bicliques and more than $10^{18}$ $(6,3)$-bicliques within. Therefore, enumeration-based counting methods including BCList++ (Yang et al., 2021), EPivoter (Ye et al., 2023) that produce the exact solution are not scalable. On the other hand, in many applications of biclique counting, an approximate count is often sufficient. In graph kernel methods (Sheshbolouki and Özsu, 2022), motifs serve as the basis for defining similarity measures between graphs in tasks like classification and anomaly detection. Since graph kernels depend on relative similarities rather than exact counts, approximate counts can preserve kernel performance while significantly reducing the computational overhead. This efficiency facilitates similarity computations in fields such as bioinformatics (e.g., comparing protein interaction networks) and cybersecurity (e.g., detecting similar attack patterns across networks). In recommendation systems (Wang et al., 2014), biclique counting helps uncover dense substructures among users or items, such as groups of users with shared interests or items commonly co-purchased. In large-scale environments such as e-commerce or streaming platforms, allowing a small error margin for counting can significantly reduce the runtime cost while still preserving the effectiveness of the downstream tasks.
@@
From this perspective, we can see that developing algorithms that produce approximate answers to $(p,q)$-biclique counting is more practical and more applicable to large-scale data. Ye et al. propose a sampling-based method called EP/Zz++ for approximate counting $(p,q)$-bicliques (Ye et al., 2023). However, its precision is not satisfactory as $p,q$ increases. For example, when querying $(5,9)$-bicliques in the data set Twitter, EP/Zz++ could produce an answer of the error ratio of more than $200\%$ [^1] Let $C$ and $\hat{C}$ be the exact and approximate counts of the $(p,q)$-biclique, respectively. The estimation error ratio is defined as $\frac{|\hat{C}-C|}{C}$. This shows the need to develop a better algorithm for approximate $(p,q)$-biclique counting, aiming to improve scalability and accuracy.
@@
In this paper, we propose a new sampling-based method that produces a high-accuracy approximate counting of $(p,q)$-bicliques. We first adapt the coloring trick in this problem, inspired by several counting methods of $k$-cliques in general graphs (Li et al., 2020; Ye et al., 2022). Then, we design a special type of subgraph that has a strong correlation with $(p,q)$-biclique, named $(p,q)$-brooms. The $(p,q)$-broom is a special type of spanning tree in a $(p,q)$-biclique, which has a fixed structure that can be taken advantage of when counting the amount. Using dynamic programming, we can count this motif efficiently and precisely. It also naturally adapts to the coloring. Finally, we develop a sampling method that takes advantage of the coloring and the intermediate result of dynamic programming. Compared to the h-zigzag pattern proposed by Ep/Zz++, our $(p,q)$-brooms have a structure that is closer to $(p,q)$-cliques, and it usually has a smaller amount in the graph. As a result, it provides a better error guarantee when sampling $(p,q)$-bicliques from them.
We empirically evaluate our algorithm against state-of-the-art exact and approximate counting algorithms on nine real-world datasets. Our contributions are summarized as follows:
@@
  • Extensive experiments show that our algorithm consistently outperforms state-of-the-art approximate algorithms, with up to 8$\times$ reduction in approximation errors and up to 50×50$\times$50 × speed-up in running time.  [^2] Code for the implementation of our method and the reproduction for all experiments can be found at https://github.com/lwn16/Biclique.
Outline. The rest of the paper is organized as follows. We review the related work in Section 2, and introduce notations and definitions in Section 3. Section 4 presents our sampling-based algorithm. Experimental results are given in Section 5, and we conclude in Section 6.
@@ 2. Related Works
In this section, we first review the existing works on biclique and $k$-clique counting problems and then briefly review other motif counting methods.
Biclique Counting. The biclique counting problem focuses on enumerating $(p,q)$-bicliques within bipartite graphs, with particular emphasis on the $(2,2)$-biclique, commonly known as a butterfly—a fundamental structural pattern with significant real-world applications. Given the importance of butterfly motifs, researchers have developed numerous algorithms for their efficient enumeration and counting (Wang et al., 2019; Sanei-Mehri et al., 2018; Wang et al., 2014). The state-of-the-art method utilizes the vertex priority and cache optimization (Wang et al., 2019). Recent advances have emerged in multiple directions: parallel computing techniques that optimize both memory and time efficiency (Xia et al., 2024), I/O-efficient methods (Wang et al., 2023) that minimize disk operations, and approximation strategies that provide accurate counts while reducing computational overhead. The field has further expanded to address more complex graph types, including uncertain graphs with probabilistic edges (Zhou et al., 2021) and temporal graphs incorporating time-varying relationships (Cai et al., 2024). The $(p,q)$-biclique counting is a general version of butterfly counting, which recently gained much attention (Ye et al., 2023; Yang et al., 2021). The state-of-the-art methods are extensively introduced and analyzed in the latter (See Section 3.1).
@@
Other Motif Counting. Beyond $k$-clique and $(p,q)$-biclique, numerous works are focusing on counting general motifs, such as four-cycles and other subgraph patterns (Hou et al., 2024; Zhao et al., 2023; Bressan et al., 2018; Ma et al., 2019). These methods can be broadly divided into two categories: traditional counting approaches (Bressan et al., 2018; Ma et al., 2019) and learning-based methods (Hou et al., 2024; Zhao et al., 2023). The former is typically based on the sampling and enumeration techniques used to obtain the approximate and exact count of the motif, respectively. In contrast, learning-based approaches offer an alternative approximate solution by leveraging machine learning techniques. Beyond this primary categorization, subgraph counting can be further classified by scope: global counting determines subgraph frequencies across the entire graph, while local counting focuses on specific nodes or edges.
@@ 3. Preliminaries
Given a bipartite graph $G=((U,V),E)$, a $(p,q)$-biclique is a subgraph $G^{\prime}=((U^{\prime},V^{\prime}),E^{\prime})\in G$ where $|U^{\prime}|=p$, $|V^{\prime}|=q$ and $E^{\prime}=\{(u,v)\mid u\in U^{\prime},v\in V^{\prime}\}$ (a complete bipartite graph).
@@ 4. ALGORITHM
In this section, we propose a new sampling-based algorithm called Colored Broom-based Sampling (CBS) that solves 1 approximately.
Inspired by methods for solving $k$-clique counting problems on general graphs (Li et al., 2020; Ye et al., 2022), we start by assigning a color to every vertex on the graph (Section 4.1). Our coloring method is simple, and produces assignments with a small number of colors empirically.
Instead of counting bicliques directly, our next step is to calculate the number of a special motif called ”broom” that we design (Section 4.2). Specifically, the $(p,q)$-broom is defined as follows:
@@
Given a bipartite graph $G((U,V),E)$, two vertex sets $U^{\prime}\in U$ and $V^{\prime}\in V$ of size $p$ and $q$, respectively. A vertex ordering is also given: $\{u_{1},u_{2},\dots,u_{p}\}$ and $\{v_{1},v_{2},\dots,v_{q}\}$. The corresponding $(p,q)$-broom is the subgraph $G^{\prime}((U^{\prime},V^{\prime}),E^{\prime})\in G$ that satisfies $E^{\prime}=\{(u_{i},v_{\lfloor\frac{(i-1)(q-1)}{p-1}+1\rfloor})\mid 1\leq i\leq p\}\cup\{(u_{\lceil\frac{(i-1)(p-1)}{q-1}\rceil},v_{i})\mid 2\leq i\leq q\}.$
@@
In Figure 2, we provide an illustration of a $(6,3)$-broom. We can observe that such a subgraph is, in fact, a tree, which contains $p+q-1$ edges and is connected. As shown in Figure 2, we can clearly see that each vertex $v_{i}$ in $V$ connects to the $2-3$ vertices $u_{j}$ in $U$ that $i$ and $j$ are close. Note that the defined ordering is only used to characterize the pattern, providing a better visualization, and it does not have to align with the original vertex index. Visually speaking, a $(p,q)$-broom is the ”skeleton” of a $(p,q)$-biclique. The general idea here is to design a way to sparsify the dense biclique while trying best to preserve its identity. This broom pattern is very different from EP/Zz++ (Ye et al., 2023)’s $h$-zigzag patterns, as they only use a simple path. We then empirically show that such pattern’s count can lead to a much better approximation of biclique counting.
There are at least two observations for the broom:
  1. (1)
    A spanning tree captures a subgraph better than a simple path.
  2. (2)
    All $p+q-1$ edges are allocated to each vertex nearly evenly.
This distribution design limits the total number of existing brooms, which makes the error analysis tight. Despite its complicated structures, we show in Section 4.2 that its amount can be computed precisely and efficiently using a dynamic programming process with respect to color assignment.
@@ 4.1. Coloring
The color assignment should guarantee that for any $(p,q)$-clique $H$, there are no two vertices in $U(H)$ that share the same color. Similarly, the same is true for any two vertices in $V(H)$.
@@
$S$ denotes the set of vertices that has not been colored. Whenever $S$ is not empty (Algorithm 1), we will try coloring some vertices with a new color, denoted as $\kappa$. For brevity, we initialize $\kappa$ to be $1$ (Algorithm 1), and whenever we need a new color, we increase $\kappa$ by $1$ (Algorithm 1). $Cnt$ is a temporary array that is used to guarantee the coloring is legal. Specifically, when we try coloring with color $\kappa$, $Cnt[u]$ stores the number of $v$ that satisfies $v\in N(u,G)$ and there exists $w\in N(v,G)$ such that $c(w)=\kappa$. In each round of coloring, we start by initializing $Cnt$ to be all zeros (Algorithm 1). We use $S_{next}$ to store vertices that have to be colored in the next round, initialized as $\emptyset$ (Algorithm 1). We enumerate vertices in $S$ in a randomized order (Algorithm 1). For each enumerated vertex $u$, we first set its color to be $\kappa$ (Algorithm 1) and update $Cnt$ correspondingly (Algorithm 1). If there exists a vertex $x$ such that $Cnt[x]\geq q$ (Algorithm 1), then there could exist a $(p,q)$-biclique with two vertices with the same color. Therefore, such coloring is illegal, and we cannot color $u$ with $\kappa$. In such case, we first revert the changes in $Cnt$ (Algorithm 1) and add $u$ into $S_{next}$ (Algorithm 1). In the end, we have tried every vertex in $S$ at least once and have deferred some vertices’ coloring to upcoming rounds. We assign $S_{next}$ to $S$ and continue (Algorithm 1).
It is not hard to see that the runtime of this coloring process is dominated by the total time of the total number of colors, i,e. $\kappa$. Note that querying, updating, and reverting $cnt$ can be done in $O(|N(u,G)|)$ each operation. In all, the time complexity is $O(\kappa|E|)$. Shown in Table 1, on real-world datasets, $\kappa$ is not large and our coloring algorithm runs very efficiently in turn. The only question left is how such an algorithm guarantees a correct color assignment. We prove it in the following lemma.
@@
After executing Coloring$(G,p,q)$, for any $(p,q)$-clique $H$ in $G$, there are no two vertices $x,y$ either both in $U(H)$ or in $V(H)$ that share the same color, i.e. $c(x)=c(y)$.

### v2

@@ Abstract.
[…]
copyright: none
PVLDB Reference Format:
PVLDB, 19(4): XXX-XXX, 2025.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 4 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/lwn16/Biclique.
@@ 1. Introduction
The bipartite graph stands as a cornerstone in graph mining, comprising two distinct sets of vertices where edges exclusively link vertices from different sets. This concept serves as a powerful tool for modeling relationships across various real-world domains, including recommendation networks (Chen et al., 2020), collaboration networks (Beutel et al., 2013), and gene coexpression networks (Kaytoue et al., 2011). For example, in recommendation networks, users and items typically constitute two distinct vertex types. The interactions between vertices form a bipartite graph, with edges representing users’ historical purchase behaviors.
@@
Despite its importance, counting $(p,q)$-bicliques is very challenging due to its exponential increase with respect to $p$ and $q$ (Yang et al., 2021). For example, in one of the graph datasets Twitter, with fewer than $2\times 10^{6}$ edges, there are more than $10^{13}$ $(5,4)$-bicliques and more than $10^{18}$ $(6,3)$-bicliques within. Therefore, enumeration-based counting methods including BCList++ (Yang et al., 2021), EPivoter (Ye et al., 2023) that produce the exact solution are not scalable. On the other hand, in many applications of biclique counting, an approximate count is often sufficient. In graph kernel methods (Sheshbolouki and Özsu, 2022), motifs serve as the basis for defining similarity measures between graphs in tasks like classification and anomaly detection. Since graph kernels depend on relative similarities rather than exact counts, approximate counts can preserve kernel performance while significantly reducing the computational overhead. This efficiency facilitates similarity computations in various fields:
  • Bioinformatics: Zhang et al. (Zhang et al., 2014) proposed algorithms to enumerate maximal bicliques in bipartite graphs derived from functional genomics (genes vs. gene-sets), enabling the identification of co-regulatory modules. A high number of bicliques in certain regions often indicates dense regulatory relationships, which can be helpful for detecting such modules. However, full enumeration quickly becomes infeasible as graphs grow larger or the number of bicliques explodes. This limitation motivates our work on scalable approximate biclique counting, which aims to estimate biclique numbers (or those of particular sizes) accurately without exhaustive enumeration.
  • Cybersecurity: In e-commerce fraud detection, bicliques also serve as key indicators of suspicious activity. For example, Wu et al. (Wu et al., 2024) show that in Taobao’s large transaction network, click-farming attacks can be modeled as large bicliques. Similarly, Ban (Ban, 2018) introduces the Maximal Half-Isolated Biclique problem to detect synchronized fraudulent behavior in customer–product graphs. These studies highlight that high biclique counts—or the presence of unusually large bicliques—in specific regions of a bipartite graph can be strong signals of fraud.
  • Data Mining: In recommendation systems (Wang et al., 2014), biclique counting helps uncover dense substructures among users or items, such as groups of users with shared interests or items commonly co-purchased. In large-scale environments such as e-commerce or streaming platforms, allowing a small error margin for counting can significantly reduce the runtime cost while still preserving the effectiveness of the downstream tasks.
@@
From this perspective, we can see that developing algorithms that produce approximate answers to $(p,q)$-biclique counting is more practical and more applicable to large-scale data. Ye et al. propose a sampling-based method called EP/Zz++ for approximate counting $(p,q)$-bicliques (Ye et al., 2023). However, its precision is not satisfactory as $p,q$ increases. For example, when querying $(5,9)$-bicliques in the data set Twitter, EP/Zz++ could produce an answer of the error ratio of more than $200\%$. [^1] Let $C$ and $\hat{C}$ be the exact and approximate counts of the $(p,q)$-biclique, respectively. The estimation error ratio is defined as $\frac{|\hat{C}-C|}{C}$. This shows the need to develop a better algorithm for approximate $(p,q)$-biclique counting, aiming to improve scalability and accuracy.
@@
In this paper, we propose a new sampling-based method that produces a high-accuracy approximate counting of $(p,q)$-bicliques. We first adapt the coloring trick in this problem, inspired by several counting methods of $k$-cliques in general graphs (Li et al., 2020; Ye et al., 2022). Then, we design a special type of subgraph that has a strong correlation with $(p,q)$-biclique, named $(p,q)$-brooms. The $(p,q)$-broom is a special type of spanning tree in a $(p,q)$-biclique, which has a fixed structure that can be taken advantage of when counting the amount. Using dynamic programming, we can count this motif efficiently and precisely. It also naturally adapts to the coloring. Finally, we develop a sampling method that takes advantage of the coloring and the intermediate result of dynamic programming. Compared to the h-zigzag pattern proposed by Ep/Zz++, our $(p,q)$-brooms have a structure that is closer to $(p,q)$-cliques, and they usually have a smaller amount in the graph. As a result, it provides a better error guarantee when sampling $(p,q)$-bicliques from them. Our algorithm is still heuristic since it includes a graph-coloring process, and the sampling time analysis contains a ratio that could be exponential. However, these processes are shown to be very efficient in practice. We empirically evaluate our algorithm against state-of-the-art exact and approximate counting algorithms on nine real-world datasets. All experimental results demonstrate the effectiveness of our algorithm. Besides the mathematical analysis, the high-level idea here is the sparsification. Since $(p,q)$-bicliques in real-world datasets contain realistic meanings, finding a sparsifier that is likely to preserve the information within bicliques is a promising approach to reducing the computational cost in practice. On the other hand, the mathematical analysis of unbiasedness and sampling time ensures that we can use this algorithm in practice. We can set the sampling time to a large number, allowing us to withstand the overall cost. Or we can iteratively increase the sampling time and terminate while the result shows a convergence trend.
Our contributions are summarized as follows:
@@
[…]
  • Additional experiments further support the effectiveness and novelty of our algorithm, which includes empirical results such as an ablation study on the newly designed coloring method and an experiment on the hyper-parameters.
Outline. The rest of the paper is organized as follows. We review the related work in Section 2, and introduce notations and definitions in Section 3. Section 4 presents our sampling-based algorithm. Experimental results are given in Section 5 and Section 6. We present major experimental results in Section 5 and provide several detailed analyses. We conclude in Section 7.
@@ 2. Related Works
Biclique Counting. The biclique counting problem focuses on enumerating $(p,q)$-bicliques within bipartite graphs, with particular emphasis on the $(2,2)$-biclique, commonly known as a butterfly—a fundamental structural pattern with significant real-world applications. Given the importance of butterfly motifs, researchers have developed numerous algorithms for their efficient enumeration and counting (Wang et al., 2019; Sanei-Mehri et al., 2018; Wang et al., 2014). The state-of-the-art method utilizes the vertex priority and cache optimization (Wang et al., 2019). Recent advances have emerged in multiple directions: parallel computing techniques that optimize both memory and time efficiency (Xia et al., 2024), I/O-efficient methods (Wang et al., 2023) that minimize disk operations, and approximation strategies that provide accurate counts while reducing computational overhead. The field has further expanded to address more complex graph types, including uncertain graphs with probabilistic edges (Zhou et al., 2021) and temporal graphs incorporating time-varying relationships (Cai et al., 2024). The $(p,q)$-biclique counting is a general version of butterfly counting, which recently gained much attention (Ye et al., 2023; Yang et al., 2021). The state-of-the-art methods are introduced and analyzed in the latter (See Section 3.1).
@@
Other Motif Counting. Beyond $k$-clique and $(p,q)$-biclique, numerous works are focusing on counting general motifs, such as four-cycles and other subgraph patterns (Hou et al., 2024; Zhao et al., 2023; Bressan et al., 2018; Ma et al., 2019; Seshadhri and Tirthapura, 2019; Zhou et al., 2024). These methods can be broadly divided into two categories: traditional counting approaches (Bressan et al., 2018; Ma et al., 2019; Seshadhri and Tirthapura, 2019) and learning-based methods (Hou et al., 2024; Zhao et al., 2023). The former is typically based on the sampling and enumeration techniques used to obtain the approximate and exact count of the motif, respectively. In contrast, learning-based approaches offer an alternative approximate solution by leveraging machine learning techniques. Beyond this primary categorization, subgraph counting can be further classified by scope: global counting determines subgraph frequencies across the entire graph, while local counting focuses on specific nodes or edges.
Dynamic Programming in Subgraph Problems In our proposed algorithm, dynamic programming is used to compute the number of special types of spanning trees. In many recent works for subgraph matching algorithms (Kim et al., 2021; Han et al., 2019), they also use dynamic programming to search a data graph for embeddings of a path tree obtained from a query graph. Although we all fix an ordering for these trees and conduct dynamic programming according, our method focuses and works differently: We use both the intermediate and final results to ensure the following sampling process can be effective. Furthermore, our transition formula and recurrence relations are significantly different. The key idea of our method is to take advantage of the bipartite structure. We split the tree into several ”prefix” subtrees and identify them by the number of edges and the last edge that adds in. In this way, we manage to compute the number of all trees as well as the number of all ”prefix” tree.
@@ 3. Preliminaries
Given a bipartite graph $G=((U,V),E)$, a $(p,q)$-biclique is a subgraph $G^{\prime}=((U^{\prime},V^{\prime}),E^{\prime})\in G$ where $|U^{\prime}|=p$, $|V^{\prime}|=q$ and $E^{\prime}=\{(u,v)\mid u\in U^{\prime},v\in V^{\prime}\}.$
@@ 4. ALGORITHM
In this section, we propose a new algorithm called Colored Broom-based Sampling (CBS) that solves 1 approximately.
1. Coloring.
Inspired by methods for solving $k$-clique counting problems on general graphs (Li et al., 2020; Ye et al., 2022), we start by assigning a color to every vertex on the graph (Section 4.1). Our coloring method is simple, and produces assignments with a small number of colors empirically. Although graph coloring has already been used in k-clique problems on general graphs, it is a new attempt to utilize it on bipartite graphs for biclique problems. On the other hand, if we directly apply the traditional coloring, it becomes 2-coloring and cannot ensure the distinct color assignment for every existing biclique. In general, coloring can be considered as an optimization aiming to reduce the number of accountable brooms. Intuitively, each $(p,q)$-biclique must obtain at least $p$ or $q$ colors due to the definition. In turn, increase the probability such that the sampled broom is a biclique. It also allows us to construct a dynamic programming implemented by following the color indices.
We construct a new coloring condition and prove its effectiveness in result. Since the graph-coloring problem is NP-Complete, there should be no theoretically optimal guarantee. In our empirical study, we have conducted an ablation study (Section 6.1) showing the effectiveness of the coloring process as well as slightly different coloring strategies.
2. Broom Pattern Counting.
Instead of counting bicliques directly, our next step is to calculate the number of a special motif called ”broom” that we design (Section 4.2). It is a special type of spanning trees of $(p,q)$-bicliques. Our first insight is that a spanning tree captures a subgraph better than a simple path. Then, since in the real-world dataset, $(p,q)$-bicliques contain realistic meanings and we also hope the sparsified motif can preserve this inner information. Therefore, we try to allocate all p+q1p+q-1 edges to each vertex as evenly as possible. In turn, we have the following mathematical definition of $(p,q)$-brooms that indicates this purpose:
@@
Given a bipartite graph $G((U,V),E)$, two vertex sets $U^{\prime}\in U$ and $V^{\prime}\in V$ of size $p$ and $q$, respectively. A vertex ordering is also given: $\{u_{1},u_{2},\dots,u_{p}\}$ and $\{v_{1},v_{2},\dots,v_{q}\}$. The corresponding $(p,q)$-broom is the subgraph $G^{\prime}((U^{\prime},V^{\prime}),E^{\prime})\in G$ that satisfies
$E^{\prime}=\{(u_{i},v_{\lfloor\frac{(i-1)(q-1)}{p-1}+1\rfloor})\mid 1\leq i\leq p\}$
$\cup\{(u_{\lceil\frac{(i-1)(p-1)}{q-1}\rceil},v_{i})\mid 2\leq i\leq q\}.$
@@
In Figure 2, we provide an illustration of a $(6,3)$-broom. We can observe that such a subgraph is, in fact, a tree, which contains p+q1p+q-1 edges and is connected. As shown in Figure 2, we can clearly see that each vertex $v_{i}$ in $V$ connects to the 232-3 vertices $u_{j}$ in $U$ that $i$ and $j$ are close. Note that the defined ordering is only used to characterize the pattern, providing a better visualization, and it does not have to align with the original vertex index. Visually speaking, a $(p,q)$-broom is the ”skeleton” of a $(p,q)$-biclique. The general idea here is to design a way to sparsify the dense biclique while trying best to preserve its identity. This broom pattern is very different from EP/Zz++ (Ye et al., 2023)’s $h$-zigzag patterns, as they only use a simple path. We then empirically show that such pattern’s count can lead to a much better approximation of biclique counting. This distribution design limits the total number of existing brooms, which makes the error analysis tight. Despite its complicated structures, we show in Section 4.2 that its amount can be computed precisely and efficiently using a dynamic programming process with respect to color assignment.
3. Counting by Sampling Brooms.
@@ 4.1. Coloring
The color assignment should guarantee that for any $(p,q)$-clique $H$, there are no two vertices in $U(H)$ that share the same color. Similarly, the same is true for any two vertices in $V(H)$. The way to achieve this is to relax the condition as ”limiting same-color-endpoint-wedges $\leq q$”, which is easy to keep track of.
@@
$S$ denotes the set of vertices that has not been colored. Whenever $S$ is not empty (Algorithm 1), we will try coloring some vertices with a new color, denoted as $\kappa$. For brevity, we initialize $\kappa$ to be $1$ (Algorithm 1), and whenever we need a new color, we increase $\kappa$ by $1$ (Algorithm 1). $Cnt$ is a temporary array that is used to guarantee the coloring is legal. Specifically, when we try coloring with color $\kappa$, $Cnt[u]$ stores the number of $v$ that satisfies $v\in N(u,G)$ and there exists $w\in N(v,G)$ such that $w\neq u$ and $c(w)=\kappa$. In each round of coloring, we start by initializing $Cnt$ to be all zeros (Algorithm 1). We use $S_{next}$ to store vertices that have to be colored in the next round, initialized as $\emptyset$ (Algorithm 1). We enumerate vertices in $S$ in a randomized order (Algorithm 1). For each enumerated vertex $u$, we first set its color to be $\kappa$ (Algorithm 1) and update $Cnt$ correspondingly (Algorithm 1). If there exists a vertex $x$ such that $Cnt[x]\geq q$ and $c(x)=\kappa$ (Algorithm 1), then there could exist a $(p,q)$-biclique with two vertices with the same color. Therefore, such coloring is illegal, and we cannot color $u$ with $\kappa$. In such case, we first revert the changes in $Cnt$ (Algorithm 1) and add $u$ into $S_{next}$ (Algorithm 1). In the end, we have tried every vertex in $S$ at least once and have deferred some vertices’ coloring to upcoming rounds. We assign $S_{next}$ to $S$ and continue (Algorithm 1).
To see that the runtime of this coloring process is dominated by the total time of the total number of colors, i,e. $\kappa$, we need to show that querying, updating, and reverting cntcnt can be done in $O(|N(u,G)|)$ each operation. Here we provide one possible implementation about how to achieve $|N(u,G)|$ for a single update as follows: The idea is to use the lazy propagation trick. We only maintain cntcnt from vertices that are (will be) colored $\kappa$, and each round we maintain a status $status[v]=(t,w)$ for $V$ during each round coloring $U$. Specifically, $t=0$ indicates $v$ has no neighbor of color $\kappa$; $t=1$ means exactly one neighbor $w$ of color $\kappa$, and $t=2$ means at least two neighbors are colored $\kappa$. Whenever we color $u$ to $\kappa$, we need to check all its neighbor $v\in N(u,G)$. Let $status[v]=(t,w)$. If $t\geq 1$,we update $Cnt[u]$. If $t=1$, we update $Cnt[w]$. After it, we update $status[v]$ accordingly. In all, for each $u$, we update its neighbor $v$ with $O(1)$ operations. Therefore, the total cost for a single update is $O(|N(u,G)|)$. In all, the time complexity is $O(\kappa|E|)$. Note that $\kappa$ is naturally bounded by $|U(H)|$. In fact, shown in Table 1, on real-world datasets, $\kappa$ is not large and our coloring algorithm runs very efficiently in turn. The only question left is how such an algorithm guarantees a correct color assignment. We prove it in the following lemma.
@@
After executing Coloring(G,p,q)(G,p,q), for any $(p,q)$-clique $H$ in $G$, there are no two vertices $x,y$ either both in $U(H)$ or in $V(H)$ that share the same color, i.e. $c(x)=c(y)$.